OpenStudy (anonymous):

what is 4x-22+34 divided by 2

OpenStudy (anonymous):

I need x's value

OpenStudy (anonymous):

4x-22+34 = 4x+12 take common ( 2) 4x+12 = 2(2x+6) = 2(2)(x+3) divide by 2 4(x+3) /2 = 2(x+3) = 2x+6